Question 1 What is the wavelength of an electromagnetic wave of frequency 98.4-MH7 Quiz navigation Not yet answered Marked out of 1.00 r Flag question Answer Finish attempt... Time left 0:19:47 Question 2 Suppose you have visible light whöse wavelength in vacuum is 450nm. If the light travels through glass, what happens to the wavelength of the light? Select one: O A Wavelength increases O B. Wavelength decreases Not yet answered Marked out of 1.00 P Flag C. Wavelength stays the same D. I dont know because I didnt read the text ome distance away, as series Question 3 two slits and then lands on a scree When light passes through
) : U TRU Moodle English (en) C.Wavelength stays the same D. I dont know because I didnt read the text Search Courses Go A Ling O O Question 3 Not yet answered Marked out of 1.00 P lag question When light passes through two slits and then lands on a screen some distance away, as series of bright and dark fringes are created. The bright fringes are found at angles given by At the second bright fringe the value of m is 2. This means the wave from one slit has travelecd Select one: O A. the same distance as the wave from the other slit O B. one wavelength farther than the wave from the other slit Ο-C. one and a half wavelengths farther than the wave from the other slit D. two wavelengths farther than the wave from the other slit O E. two and a half wavelengths farther than the wave from the other slit White light alls n a difaction grating creating a white central peak. As you move out from this central peak Question 4 Not yet answered what is the first visible color you encounter? Marked out of 1.00 Select one: A. red
NTETRU Moodle English (en) Search Courses Go Lingging Lua Question 4 Not yet answered Marked out of 1.00 White light falls on a diffraction grating creating a white central peak. As you move out from this central pealk what is the first visible color you encounter? Select one: A, red Fag question o B. orange C, yellow o D. green O E. blue . F. violet Question 5 Visible light can uridergo interference when it reflects from or is transmitted through a thin film. What is a typically thickness for these films? Not yet answered Select one: 1.00 a. less than 1nm b. 100 to 1000 nm c.1pm to Sum Y Flag
Answer #1

1.f=98.4MHz=98.4x106Hz

the electromagnetic wave travells at the speed of light(c)

c=3x108m/sec

c=wavelength x frequency

Wavelength=c/f=3x108/98.4x106=3.0487=3.05meterAns

2.Acoording to Snell's law-

sin\theta1/sin\theta2=v​​​​​​1​​​​​/v​​​2 =n​​2/n​​​​1

where v​​​​​​1 is velocity of wave in vacuum.

v2 is velocity of wave in glass.

n1 is refractive index of vacuum=1

n2 is refractive index of glass=n>1

c/v2=n/1

v2=c/n

frequency of light is same in different mediums.

V =\lambdaf

\lambda2f=\lambda1f/n

\lambda2=\lambda1/n​​​​

n>1

so \lambda2<\lambda​​​​​​1

so wavelength of light is decreased in glass.Ans

​​​​​​